The Green Company processes unprocessed goat milk up to the split-off point where two products, condensed goat milk and skim goat milk result. The following information was collected for the month of October:

        Direct Materials processed:       105,000 gallons (after shrinkage)

        Production:                                   Condensed goat milk      45,500 gallons

                                                                Skim goat milk                  59,500 gallons

        Sales:                                              Condensed goat milk      $4.50 per gallon

                                                                Skim goat milk                  $4.00 per gallon

The costs of purchasing the of unprocessed goat milk and processing it up to the split-off point to yield a total of 105,000 gallons of saleable product was $191,480. There were no inventory balances of either product. Condensed goat milk may be processed further to yield 45,000 gallons (the remainder is shrinkage) of a medicinal milk product, Xyla, for an additional processing cost of $5 per usable gallon. Xyla can be sold for $20 per gallon.

Skim goat milk can be processed further to yield 58,200 gallons of skim goat ice cream, for an additional processing cost per usable gallon of $5. The product can be sold for $12 per gallon.

There are no beginning and ending inventory balances.

Using the sales value at split-off method, what is the gross-margin percentage for skim goat milk at the split-off point? (Round intermediary percentages to the nearest hundredth.)

A.

46.25%

B.

50.00%

C.

56.75%

D.

53.75%
Answer #1

The correct answer is C) 56.75 %

Explanation and calculation

Firstly we have to calculate the total sale value at split off point

= sale value of condensed goat milk + skimmed goat milk

= (45500*4.5)+ (4*59500)

= 204750 +238000

= $ 442750

Joint cost of skimmed goat milk at split off point

= sale value of skimmed goat milk at split off / total sale value

= 238000/442750*191480

= $ 102930

Gross margin percentage = gross margin / sale value at split off point

= ((238000-102930)/238000)*100

= 56.75 %

Thus the final answer is C) 56.75 %